4. A fair two-sided coin is tossed repeatedly. (a) Find the expected number of tails until the first head is flipped. (b) Find the probability that there are exactly 5 heads in the first 10 flips. (c) Use the central limit theorem/normal approximation to approximate the probability that in the first 100 flips, between 45 and 55 of the flips are heads.
Answer #1

a)

here probability of a head for a fair coin p=1/2

therefore expected number of tails untill first head =1/p-1=1/(1/2)-1=1

b)

P(5 heads in first 10 flips)=10C5(1/2)5(1/2)5 =0.2461

c)

n= 100 p= 0.5000
here mean of distribution=μ=np= 50
and standard deviation σ=sqrt(np(1-p))= 5.0000

therefore from normal approximation of binomial distribution:

probability = P(45<X<55) = P(-1<Z<1)= 0.8413-0.1587= 0.6826